Home CMS Production Clean room MedAustron HEPHY testbeams old
testbeam-RD50-DESY-Jul2023 testbeam-RD50-CERN-Oct2022 testbeam-DESY-Oct2017 testbeam-DESY-April2017 testbeam-SPS2015 testbeam-SPS2014 testbeam-DESY14 testbeam-SPS12 testbeam-SPS11 testbeam-SPS10 testbeam-SPS09 testbeam-SPS08 LP-TPC
  HEPHY logbook for testbeam at SPS June 2009  Not logged in ELOG logo
Message ID: 59     Entry time: Sat Aug 22 01:06:34 2009
Author: Christian Irmler 
Project: BELLE 
Subject: APVDAQ + telescope 40 MHz transparent mode (cooling?) 
Run Number: run013 
Events: 75000 
StartTime: 21.08.2009 23:49:38  
EndTime: 22.08.2009 03:21:48 

Belle DUT, tranparent mode
HV=80V, 40 MHz, beam on, 120 GeV
6 samples
chiller set to 13 deg.

beam position:  p-side: APV2 of Origami, n-side: between APV1 and 2 of Origami (center of sensors), beam spot: 9x9 mm
JP-module is hit at APV3, module seems to be on a higher position than in past beam test (eventuelly flipped)

APVDAQ and telescope runs are not synchronized

trigger rate: about 300 - 500 ev / spill

there was probably no water flow throught the pipe of the origami module during this run --> must be repeated

ELOG V3.1.5-fc6679b